I guess this is similar to one I just asked, but I do not get how to approach this:

Find the midpoint of the line segment having the given endpoints.
9) (8, -4) and (-7, 8)

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(75887)

Find the midpoint of the line segment having the given endpoints.
(8, -4) and (-7, 8)
---
If (x,y) is the midpoint:
x = (-7+8)/2 = 1/2
y = (-4+8)/2 = 2
---
midpoint: (1/2 , 2)
======================
cheers,
Stan H.

Answer by solver91311(24713)



Use the mid-point formulae:

and

where and are the coordinates of the given points.

You can do your own arithmetic.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism